(1-x^2)yp^2+x^2q=0


1
Expert's answer
2021-03-07T16:50:52-0500
"(1-x^2)yp^2+x^2q=0\\
\frac{(1-x^2)yp^2}{p^2q}+\frac{x^2q}{p^2q}=\frac{0}{p^2q}\\
\frac{(1-x^2)y}{q}+\frac{x^2}{p^2}=0\\
-\frac{(x^2-1)y}{q}+\frac{x^2}{p^2}=0\\
\frac{x^2}{p^2}=\frac{(x^2-1)y}{q}=a \qquad \text{(say)}\\"

Then:


"\frac{(x^2-1)y}{q}=a \implies q= \frac{(x^2-1)y}{a}\\
\frac{x^2}{p^2}=a \implies p = \frac{x}{\sqrt{a}}"

Since


"dz = \dfrac{\partial z}{\partial x}dx + \dfrac{\partial z}{\partial y}dy"

Therefore,


"dz = \frac{x}{\sqrt{a}}dx+\frac{(x^2-1)y}{a}dy"

Integrating through:


"z = \frac{x^2}{2\sqrt{a}}+\frac{(x^2-1)y^2}{2a}+c\\
\therefore 2z = \frac{x^2}{\sqrt{a}}+\frac{(x^2-1)y^2}{a}+k\\"

where "k = 2c"